# A solution of the problem of standard compact Clifford-Klein forms

### v1

@@ 5.1.4 Completion of proof for $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ simple
[…]
  3. 3.
[…]
Proof.
Recall that by Proposition 5 the inclusions of a simple Lie algebra $\mathfrak{h}^{c}$ into classical $\mathfrak{g}^{c}$ fall into classes:
  1. 1.
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$, where $\hat{\mathfrak{h}}^{c}$ is maximal parabolic,
  2. 2.
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$, where
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}=\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k,\mathbb{C}),$
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}=\mathfrak{sp}(k,\mathbb{C}\oplus\mathfrak{sp}(n-k,\mathbb{C}).$
  3. 3.
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}$, where $\hat{\mathfrak{h}}^{c}$ is of tensor type,
  4. 4.
    $\mathfrak{h}^{c}$ is irreducible and simple.
We begin the proof with case (4) which requires the longest calculations. Since $\hat{\mathfrak{h}}^{c}$ is simple and irreducible, we consider first the case when the inclusion $\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$ is a fundamental representation. Looking at Tables 4,5,6 we calculate the possible ranks of the inclusions, using formulas from Table 4. In any case we obtain the required inequality. If the representation $\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$ is not fundamental, we use the Weyl dimension formula [18], Theorem 5.84. The proof of case $(2)$ can be obtained from the following observation. If $\mathfrak{h}\hookrightarrow\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$, then necessarily there is an embedding of $\mathfrak{h}$ to at least one of $\mathfrak{g}_{i}$ ($i=1,2$). Otherwise, the kernels of the projections onto $\mathfrak{g}_{i}$ composed with the embedding would yield non-trivial ideals in $\mathfrak{h}$. Thus, $\mathfrak{h}^{c}\subset\mathfrak{g}_{1}$ or $\mathfrak{h}^{c}\subset\mathfrak{g}_{2}$. In any case, one obtains an embedding of simple Lie algebras with $\mathfrak{g}_{i}$ of lower ranks. One proceeds by induction using $(4)$. This argument also settles case $(1)$. Since $\hat{\mathfrak{h}}^{c}$ is maximal parabolic, its reductive Levi subalgebra has the form
$\mathfrak{f}(\mathfrak{c},[M]),\,M\subset\Pi,$
where $M$ is a subset of $\Pi$, and $[M]$ is a subsystem of the root system $\Delta$. Thus, $\mathfrak{h}^{c}$ is a simple subalgebra contained in a semisimple subalgebra of a lower rank. In the same manner one settlles (3), since $\hat{\mathfrak{h}}^{c}$ is a direct sum of two simple Lie subalgebras of $\mathfrak{sl}(V_{1})\oplus\mathfrak{sl}(V_{2})$ embedded into $\mathfrak{sl}(V_{1})\otimes\mathfrak{sl}(V_{2})$. Thus, the same argument works.
Corollary 6.
If $\mathfrak{h}^{c}$ belongs to any of the inclusions in Proposition 7 then no real form of the pair $(\mathfrak{g}^{c},\mathfrak{h}^{c})$ cannot generate compact standard Clifford-Klein form.
Proof.
The proof follows from the proof of Proposition 6 with small modifications. One can notice that if $\mathfrak{g}^{c}=D_{n}$, or $\mathfrak{g}^{c}=B_{n}$ then the same argument works, this time applied to $\mathfrak{h}^{c}$ and the ambient Lie algebra $\mathfrak{g}^{c}$. Still, one obtains the embeddings
$\mathfrak{so}(q,2p-1-q)\subset\mathfrak{so}(m,2n+1-m),\,\,\text{or}\,\,\mathfrak{so}(q,2p-1-q)\subset\mathfrak{so}(m,2n-m)$
and the same argument using Theorem 12 completes the proof. The case $\mathfrak{g}^{c}=C_{n}$ need not be considered for the dimensional reasons: the smallest possible dimension of a symplectic representation of $D_{p}$ is $\frac{n}{2}$. ∎
Let us now list all simple real forms of non-compact type. By a slight abuse of notation we will treat simple but not absolutely simple real Lie algebras of rank 2n as Lie algebras of type $n.$ For instance $\mathfrak{g}=\mathfrak{sl}(n+1,\mathbb{C})$ is classified as type $A_{n}.$ Notice that in this case $\mathfrak{g}^{c}=\mathfrak{sl}(n+1,\mathbb{C})\oplus\mathfrak{sl}(n+1,\mathbb{C}).$
@@
Remarks:
@@
    For all classical types with $\textrm{rank}=t$ and $\textrm{rank}_{\mathbb{R}}=a$ one of the following three algebras has the highest $d$
@@
    For $\mathfrak{g}_{2}(\mathbb{C})$ one can find a real form of type I ($\mathfrak{so}(2,7)$) with the same real rank, with greater or equal $d$ and with rank at most half of the rank of the representation into algebra of rank 8 or greater. As the simplest representation of $\mathfrak{g}_{2}(\mathbb{C})$ is $B_{6}.$ Thus in each case one has to verify separately the embeddings of $\mathfrak{g}_{2}(\mathbb{C})$ into lie algebras of rank $6$ and $7$
@@
    For $\mathfrak{g}_{2(2)}$ one can find a real form of type I ($\mathfrak{so}(2,5)$) with the same real rank, with greater or equal $d$ and with rank at most half of the rank of the representation into algebra of rank 6 or greater. As the simplest representation of $\mathfrak{g}_{2(2)}$ is $B_{3}$ then one should in each case verify the embeddings of $\mathfrak{g}_{2(2)}$ into lie algebras of rank $4$ and $5$ This is because the real rank of $\mathfrak{g}_{2(2)}$ equals 2 and so if $(\mathfrak{g},\mathfrak{g}_{2(2)},\mathfrak{l})$ gives a standard compact Clifford-Klein form then necessarily the real rank of $\mathfrak{g}$ is at lest 3. If the $\textrm{rank}(\mathfrak{g})=3$ then $\mathfrak{g}$ is split and we can use Theorem “split”.
@@
Proposition 8.
For any classical absolutely simple $\mathfrak{g}$ in every case described above
$D:=d(\mathfrak{g})-d(\mathfrak{h})-d(\mathfrak{l})>0$
and so one cannot obtain compact standard Clifford-Klein forms.
Proof.
By the remarks in the previous Section we only have to check the inequality for types I,II,III (and $G_{2}$ for low ranks).
By $X_{t}$ we will denote an arbitrary type of rank $t.$ By $(i,j),$ $i,j\in\{I,II,III\}$ we denote a pair of subalgebras of type i and j, respectively. We exclude the cases in which $\mathfrak{g}$ is split, becuse of [6].
Type: $A_{2n}$ Case $\mathfrak{g}=\mathfrak{su}(a+b,2n+1-a-b)$ We start with $\mathfrak{h}$ of type $G_{2}.$ First we have to check $\mathfrak{g}_{2(2)}$ in $A_{4}.$ But in this case real rank of $\mathfrak{g}$ is at most 2. Secondly we have $\mathfrak{g}_{2}(\mathbb{C})$ in $A_{6}.$ So necessarily $\mathfrak{g}=\mathfrak{su}(3,4).$ One easily checks that there is no simple $\mathfrak{l}$ of real rank one with $d(\mathfrak{l})>8.$
@@
If $\mathfrak{h}$ is simple but not absolutely simple then it is obviously of type $X_{p}$ with $p\leq n.$ Thus we have to check Lie subalgebras of type I,II and III for $t=n.$ We have
@@
Type: $A_{2n+1}$
Again, by Proposition 7 we see that if $X_{p}\subset A_{2n+1}$ is classical simple non-regular subalgebra or simple but not absolutely simple subalgebra then $p\leq n+1.$
@@
Type: $C_{n}$
In this type we only have to check $\mathfrak{g}=\mathfrak{sp}(a+b,n-a-b).$
We begin with subalgebras of type $G_{2}.$ We can eliminate (in the case $\mathfrak{h}=\mathfrak{g}_{2(2)}$) $C_{4}$ and $C_{5}$ since the real of $\mathfrak{sp}(2,2)$ and $\mathfrak{sp}(2,3)$ equals $2$ If $\mathfrak{h}:=\mathfrak{g}_{2}(\mathbb{C})\subset A_{6}$ then (since the real rank of $\mathfrak{g}$ has to be greater than 2 and we omit the split case) necessarily $\mathfrak{g}=\mathfrak{sp}(3,3).$ But now $\mathfrak{l}\subset\mathfrak{sp}(3,3)$ has to have real rank equal to 1 and $d=22.$ One easily verifies that this is impossible. We can use the same argument for $C_{7}$ (in this case necessarily $\mathfrak{g}=\mathfrak{sp}(3,4)$).
Regular proper subalgebras of $C_{n}$ are given by $A_{p},C_{p},D_{s},$ $p+1,s\leq n.$ By Proposition 7 to obtain non-regular subalgebra in $C_{n}$ of rank greater than $\frac{n}{2}$ one has to “use” the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $C_{n}.$ By Corollary 6 any non-regular simple subalgebra in $D_{p}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
@@
We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then $\mathfrak{g}=\mathfrak{so}(3,6),$ $\mathfrak{so}(4,5).$ If $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then $\mathfrak{g}=\mathfrak{so}(s,13-s),\mathfrak{so}(s,15-s),$ $s=3,4,5,6,7.$ One verifies that there does not exist $\mathfrak{l}$ with adequate $r,d$ with one exception:
@@
Regular proper subalgebras of $B_{n}$ are given by $A_{p},B_{p},D_{s},$ $p+1,s\leq n.$ By Corollary 6 to obtain non-regular subalgebra in $B_{n}$ of rank greater than $\frac{n}{2}$ one has to consider the chain
@@
Type: $D_{n}$
Notice that if $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then we have to check $D_{6}$ and $D_{7}.$ But if $\mathfrak{g}_{2}(\mathbb{C})$ is contained in Lie algebra of type $D_{p}$ then necessarily $p\geq 8.$
Regular proper subalgebras of $D_{n}$ are given by $A_{p},D_{p},$ $p\leq n.$ Corollary 6 eliminates the possibility $\operatorname{rank}\mathfrak{h}^{c}>\frac{1}{2}\operatorname{rank}\mathfrak{g}^{c}$. The opposite inequality is analyzed below.
@@
By Corollary 6 any non-regular simple subalgebra in $B_{n}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
@@
We see that one only needs to check I and III. Also $n\geq 4,$ $\frac{n-1}{2}\leq a+b\leq\frac{n}{2}$ and so $a^{2}+b^{2}\leq\frac{n^{2}}{4}-2ab.$ (I,I) $D=n^{2}-n-2an+a+a^{2}-2bn+b+b^{2}\geq-n+a^{2}+a+b^{2}+b>0,$ unless $a=b=1$ and $n=4.$ But in this case we have embedding $B_{3}\subset D_{4},$ which is symmetric and we can conclude by [32].
@@
We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then we have to check $\mathfrak{g}=\mathfrak{so}(3,5),$ $\mathfrak{so}(3,7),$ $\mathfrak{so}(4,6).$ One verifies that there does not exist $\mathfrak{l}$ with adequate $r,d.$
We have the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $D_{n}.$ By Corollary 6 we can eliminate this case. Thus we may assume that each subalgebra of $\mathfrak{so}(a+b,2n-a-b)$ is of rank at most $\frac{n}{2}.$
@@
Proposition 9.
No real form of the triple $(\mathfrak{g}^{c},\mathfrak{h}^{c},\mathfrak{l}^{c})$ with at least one of the pair $(\mathfrak{h}^{c},\mathfrak{l}^{c})$ satisfies the assumptions 1) or 3) of Proposition 5 can yield a standard compact Clifford-Klein form.
Proof.
Consider first $\mathfrak{h}^{c}$ or $\mathfrak{l}^{c}$ which are simple, reducible and contained in direct sums of simple classical Lie algebras $\mathfrak{g}^{c}=\mathfrak{g}_{1}^{c}\oplus\mathfrak{g}^{c}_{2}$. One can notice that such subalgebra is necessarily isomorphic to a subalgebra in one of the summands, $\mathfrak{g}_{1}^{c}$ or $\mathfrak{g}_{2}^{c}$. But since $\dim\mathfrak{g}_{1}^{c}<\dim\mathfrak{g}^{c}$ and $\dim\mathfrak{g}^{c}_{2}<\dim\mathfrak{g}^{c}$, one can apply the previous considerations to the inclusion of reducible $\tilde{\mathfrak{h}}^{c}$ or $\tilde{\mathfrak{l}}^{c}$ (isomorphic to $\mathfrak{h}^{c}$ or to $\mathfrak{l}^{c}$) into one of the $\mathfrak{g}_{1}^{c}$ or $\mathfrak{g}_{2}^{c}$. Since the non-compact dimension is the invariant of a Lie algebra itself we get the same inequality for non-compact dimensions of $\mathfrak{g},\mathfrak{h},\mathfrak{l}$ ensuring that these triples cannot generate compact standard Clifford-Klein forms.
5.1.5 Proof of Theorem 2 for the semisimple case
Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ corresponds to a standard triple and let $\mathfrak{h}=\mathfrak{h}_{1}\oplus\mathfrak{h}_{2}.$ We may also assume that $\mathfrak{g}$ is of classical type and non-split. Let
$\mathfrak{h}_{i}=\mathfrak{k}_{i}\oplus\mathfrak{a}_{i}\oplus\mathfrak{n}_{i},\ \ i=1,2$
be the compatible Iwasawa decompositions. We see that $\mathfrak{h}_{1}$ is in the centralizer of $\mathfrak{a}_{2}$ in $\mathfrak{g}.$ Similarly, $\mathfrak{h}_{2}$ is in the centralizer of $\mathfrak{a}_{1}.$ Therefore, by Theorem 4 all the roots of $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ are contained on two planes: one orthogonal to $\mathfrak{a}_{1},$ the second orthogonal to $\mathfrak{a}_{2}.$ But now, by Lemma 5 for $\Sigma,$ there exists $\beta\in\Delta_{0}^{\pm}$ such that the restriction of $\beta$ to $\mathfrak{a}_{i},$ $i=1,2,$ is non-zero.
By Lemma 7 every root of $\Delta_{0}^{\pm}$ is a sum of some two roots in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ Thus $\beta=\alpha_{1}+\sigma(\alpha_{2}),$ so that $\alpha_{1}|_{\mathfrak{a}_{2}}=0$ and $\sigma(\alpha_{2})|_{\mathfrak{a}_{1}}=0.$ It follows that $\alpha_{1}$ and $\sigma(\alpha_{2})$ are orthogonal and so the root subsystem generated by those roots is necessarily of type $B_{2}.$ Thus $\mathfrak{g}$ has to by of type $B_{t}$ or $C_{t}.$ We thus have only two possibilities (we may assume that $\alpha_{1}(iX)>0$).
Case 1. $\mathfrak{g}=\mathfrak{so}(n,m),$ $n+m$ is odd. In this case for any root $\alpha$ in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ the root $\sigma(\alpha)$ has the same projection on $\mathfrak{a}.$ But in $\mathfrak{so}(n,m)$ any long root of $\Delta$ can correspond only to a 1-dimensional root space of $\mathfrak{g}.$ It follows that $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ consists only of short roots. But now $\alpha_{1}+\alpha_{2}$ is a root because a sum of any two (non-colinear) short roots in $B_{t}$ is a root.
By Theorem 4 we see that
$S_{h}^{1}=x_{\alpha_{1}}+\sum_{l=1}^{k}a^{i}_{k+l}x_{\alpha_{k+l}}+Q_{1},$
$S_{h}^{2}=x_{\alpha_{2}}+\sum_{l=1}^{k}a^{i}_{k+l}x_{\alpha_{k+l}}+Q_{2},$
belong to $\mathfrak{h}^{c}.$ We also see that $[x_{\alpha_{1}},x_{\alpha_{2}}]$ is non-zero and it is a unique factor in $[S_{h}^{1},S_{h}^{2}]$ with maximal value on $ad_{iX}.$ Thus $0\neq Y:=[S_{h}^{1},S_{h}^{2}]\in\mathfrak{h}^{c}$ and
$ad_{\mathfrak{a}_{i}}(Y)\neq 0,\ \ i=1,2.$
But this shows that $\mathfrak{h}\neq\mathfrak{h}_{1}\oplus\mathfrak{h}_{2},$ a contradiction.
Case 2 One can deal with case 2 only by using general structure theory of Lie algebras. Indeed, let $\mathfrak{g}=\mathfrak{sp}(n,m),$ where $n\leq m.$ Since $\mathfrak{g}^{c}$ is of type $C_{p}$ therefore we may represent $\Delta$ as
$\Delta=\{\pm e_{i}\pm e_{j},\pm 2e_{i}\ |\ i\neq j,1\leq i,j\leq p\},$
where $(e_{i})_{i=1}^{p}$ is a standard basis of $\mathbb{R}^{p}.$ From the general consideration in Section 5.1.5 we know that there must exist roots $\alpha_{1}$ and $\sigma(\alpha_{2})$ in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that
  1. 1.
    $\alpha_{1}$ and $\sigma(\alpha_{2})$ are orthogonal,
  2. 2.
    $\alpha_{1}(iX)=-\sigma(\alpha_{2})(iX)\neq 0,$
  3. 3.
    $\beta=\alpha_{1}+\sigma(\alpha_{2})$ is a (long) root of $\Delta$ contained in $\Delta_{0}^{\pm},$
  4. 4.
    $\alpha_{1}|_{\mathfrak{a}_{2}}=0$ and $\sigma(\alpha_{2})|_{\mathfrak{a}_{1}}=0.$
Therefore necessarily
$\alpha_{1}=e_{i}+e_{j},\ \ \ \sigma(\alpha_{2})=e_{i}-e_{j},$
for some $i\neq j$ and $e_{i}(iX)=0,$ $e_{j}(iX)>0.$ Consider the root
$\tilde{\beta}:=\alpha_{1}-\sigma(\alpha_{2})=2e_{j}.$
We see that
$\tilde{\beta}(iX)\neq 0,\ \ \tilde{\beta}|_{\mathfrak{a}_{2}}=\sigma(\alpha_{2}),\ \ \tilde{\beta}|_{\mathfrak{a}_{1}}=\alpha_{1}.$
Thus $\tilde{\beta}$ is in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ and is non-zero on $\mathfrak{a}_{1}$ and on $\mathfrak{a}_{2}.$ But, as is shown in Section 5.1.5, the roots of $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ are contained in two planes: one orthogonal to $\mathfrak{a}_{1},$ the second orthogonal to $\mathfrak{a}_{2}.$ A contradiction. $\square$.
Remark 4.
The case of $\mathfrak{g}=\mathfrak{sp}(n,m),$ $n\leq m.$ can be settled also by case-by-case analysis as in Section 5.1.4. One shows that the only possibility is that $\mathfrak{h},\mathfrak{l}$ are of the form
$\mathfrak{h}=\oplus_{s=1}^{w_{h}}\mathfrak{sp}(n_{s},m_{s}),\ \ \ \mathfrak{l}=\oplus_{s=1}^{w_{l}}\mathfrak{sp}(n_{s},m_{s}).$
By [21] (Theorem 1.5), for any $s$ we have $m_{s}=m$ (otherwise one could construct a reductive subalgebra $\mathfrak{g}^{\prime}$ in $\mathfrak{g}$ with $d(\mathfrak{g}^{\prime})>d(\mathfrak{g})$ and $\mathfrak{a}_{g^{\prime}}=\mathfrak{a}_{h}$, which is impossible. But now we see that $\mathfrak{h}$ and $\mathfrak{l}$ have to be simple. A contradiction.
6 Addendum: tables and explicit calculations
6.1 Tables

### v3

@@ 5.3 Steps of proof of Theorem 2
    looking at Tables 4,5,3, 2 and using relations between real and complex ranks together with the constraint given by inequality (6), one obtains the desired inequality for the non-compact dimensions.
As we have mentioned, the details of the calculation are postponed until Subsection 5.3.1.
Clearly, Proposition 4 exhaust all the possibilities for non-regular triples $(\mathfrak{g}^{c},\mathfrak{h}^{c},\mathfrak{l}^{c})$ of simple complex Lie algebras and their real forms, which completes the proof of Theorem 2.
5.3.1 Proof of Proposition 4
We need to list all simple real forms of non-compact type from [33], Table 4, in the form suitable for the calculations of the non-compact dimensions. By a slight abuse of notation we will treat simple but not absolutely simple real Lie algebras of rank 2n as Lie algebras of type $n.$ For instance $\mathfrak{g}=\mathfrak{sl}(n+1,\mathbb{C})$ is classified as type $A_{n}.$ Notice that in this case $\mathfrak{g}^{c}=\mathfrak{sl}(n+1,\mathbb{C})\oplus\mathfrak{sl}(n+1,\mathbb{C}).$
@@ 5.3.1 Proof of Proposition 4
Now we make some extra analysis of this list.
Observations based on the list above
@@
    For all classical types with $\textrm{rank}=t$ and $\textrm{rank}_{\mathbb{R}}=a$ one of the following three algebras has the highest number $d:=d(\mathfrak{g})$
@@
    For $\mathfrak{g}_{2}(\mathbb{C})$ one can find a real form of type I ($\mathfrak{so}(2,7)$) with the same real rank, and $d$ the same or greater, and with rank at most half of the rank of the representation into an algebra of rank $8$ or greater. As the simplest representation of $\mathfrak{g}_{2}(\mathbb{C})$ is $B_{6},$ in each case one has to verify separately the embeddings of $\mathfrak{g}_{2}(\mathbb{C})$ into some Lie algebras of rank $6$ and $7$
@@
    For $\mathfrak{g}_{2(2)}$ one can find a real form of type I ($\mathfrak{so}(2,5)$) with the same real rank, with $d$ the same or greater, and with rank at most half of the rank of the representation into a Lie algebra of rank 6 or greater. As the simplest representation of $\mathfrak{g}_{2(2)}$ is $B_{3}$, one should verify the embeddings of $\mathfrak{g}_{2(2)}$ into Lie algebras of rank $4$ and $5$ This is because the real rank of $\mathfrak{g}_{2(2)}$ equals 2 and so if $(\mathfrak{g},\mathfrak{g}_{2(2)},\mathfrak{l})$ gives a standard compact Clifford-Klein form then necessarily the real rank of $\mathfrak{g}$ is at least 3. If the $\textrm{rank}(\mathfrak{g})=3$ then $\mathfrak{g}$ is split and we can refer to Corollary 1.
@@
Final check up of the inequality $d(\mathfrak{h})+d(\mathfrak{l})<d(\mathfrak{g})$
By the observations above we only have to check the inequality for types I,II,III (and $G_{2}$ for low ranks).
By $X_{t}$ we will denote an arbitrary type of rank $t.$ By $(i,j),$ $i,j\in\{I,II,III\}$ we denote a pair of subalgebras of type i and j, respectively. We exclude the cases in which $\mathfrak{g}$ is split, because of [8].
Type $A_{2n}$ Case $\mathfrak{g}=\mathfrak{su}(a+b,2n+1-a-b)$ We begin with $\mathfrak{h}$ of type $G_{2}.$ First we have to check $\mathfrak{g}_{2(2)}$ in $A_{4}.$ But in this case the real rank of $\mathfrak{g}$ is at most 2. Secondly we have $\mathfrak{g}_{2}(\mathbb{C})$ in $A_{6}.$ So necessarily $\mathfrak{g}=\mathfrak{su}(3,4).$ One easily checks that there is no simple $\mathfrak{l}$ of real rank one with $d(\mathfrak{l})>8.$
@@
If $\mathfrak{h}$ is simple but not absolutely simple then it is obviously of type $X_{p}$ with $p\leq n.$ Thus we have to check the Lie subalgebras of type I,II and III for $t=n.$ We have
@@
Type $A_{2n+1}$
Again, by Proposition 4 we see that if $X_{p}\subset A_{2n+1}$ is a classical simple non-regular subalgebra or simple but not absolutely simple subalgebra then $p\leq n+1.$
@@
Type $C_{n}$
In this case we only have to check $\mathfrak{g}=\mathfrak{sp}(a+b,n-a-b).$
We begin with subalgebras of type $G_{2}.$ We can eliminate (in the case $\mathfrak{h}=\mathfrak{g}_{2(2)}$) $C_{4}$ and $C_{5}$ since the real rank of $\mathfrak{sp}(2,2)$ and $\mathfrak{sp}(2,3)$ equals $2$ If $\mathfrak{h}:=\mathfrak{g}_{2}(\mathbb{C})\subset A_{6}$ then (since the real rank of $\mathfrak{g}$ has to be greater than 2 and we omit the split case) necessarily $\mathfrak{g}=\mathfrak{sp}(3,3).$ But now $\mathfrak{l}\subset\mathfrak{sp}(3,3)$ has to have real rank equal to 1 and $d=22.$ One easily verifies that this is impossible. We can use the same argument for $C_{7}$ (in this case necessarily $\mathfrak{g}=\mathfrak{sp}(3,4)$).
Regular proper subalgebras of $C_{n}$ are given by $A_{p},C_{p},D_{s},$ $p+1,s\leq n.$ By Proposition 4 to obtain non-regular subalgebra in $C_{n}$ of rank greater than $\frac{n}{2}$ one has to “use” the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $C_{n}.$ By Proposition 4 any non-regular simple subalgebra in $D_{p}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
@@
We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then $\mathfrak{g}=\mathfrak{so}(3,6),$ $\mathfrak{so}(4,5).$ If $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then $\mathfrak{g}=\mathfrak{so}(s,13-s),\mathfrak{so}(s,15-s),$ $s=3,4,5,6,7.$ One verifies that there does not exist $\mathfrak{l}$ with appropriate $r,d$, with one exception:
@@
Regular proper subalgebras of $B_{n}$ are given by $A_{p},B_{p},D_{s},$ $p+1,s\leq n.$ By Proposition 4 to obtain non-regular subalgebra in $B_{n}$ of rank greater than $\frac{n}{2}$ one has to consider the chain
@@
Type $D_{n}$
Notice that if $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then we have to check $D_{6}$ and $D_{7}.$ But if $\mathfrak{g}_{2}(\mathbb{C})$ is contained in the Lie algebra of type $D_{p}$ then necessarily $p\geq 8.$
Regular proper subalgebras of $D_{n}$ are given by $A_{p},D_{p},$ $p\leq n.$ Proposition 4 eliminates the possibility $\operatorname{rank}\mathfrak{h}^{c}>\frac{1}{2}\operatorname{rank}\mathfrak{g}^{c}$. The opposite inequality is analyzed below.
@@
By Proposition 4 any non-regular simple subalgebra in $B_{n}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
@@
We see that one only needs to check I and III. Also $n\geq 4,$ $\frac{n-1}{2}\leq a+b\leq\frac{n}{2}$ and so $a^{2}+b^{2}\leq\frac{n^{2}}{4}-2ab.$ (I,I) $D=n^{2}-n-2an+a+a^{2}-2bn+b+b^{2}\geq-n+a^{2}+a+b^{2}+b>0,$ unless $a=b=1$ and $n=4.$ But in this case we have embedding $B_{3}\subset D_{4},$ which is symmetric and we can apply [35].
@@
We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then we have to check $\mathfrak{g}=\mathfrak{so}(3,5),$ $\mathfrak{so}(3,7),$ $\mathfrak{so}(4,6).$ One verifies that there does not exist $\mathfrak{l}$ with appropriate $r,d.$
We have the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $D_{n}.$ By Proposition 4 we can eliminate this case. Thus we may assume that each subalgebra of $\mathfrak{so}(a+b,2n-a-b)$ is of rank at most $\frac{n}{2}.$
@@
6 Tables
Non-compact dimensions
[…]
$A_{t}$ $2a(t+1)-a$ $1\leq a\leq\frac{t}{2}$
$A_{t}$ $2a^{2}$ $t=2a-1$ $a\geq 2$
$B_{t}$ $a(2t+1-a)$ $1\leq a\leq t$
$C_{t}$ $4a(t-a)$ $1\leq a\leq\frac{t}{2}$
$D_{t}$ $a(2t-a)$ $1\leq a\leq t$
$D_{t}$ $2a(2a-1)$ $t=2a$
$D_{t}$ $2a(2a+1)$ $t=2a+1$
Table 4: